OpenStudy (anonymous):

If two random variables have the joint density f(x1,x2)=x1x2 for 0

OpenStudy (tkhunny):

Have you considered drawing the region and cutting it in half both horizontally and vertically - and then pondering the geometry?

OpenStudy (anonymous):

I guess that I don't know how I would draw the region?

OpenStudy (tkhunny):

We get to use integration, right?? Check to see that you have it: \(\int\limits_{0}^{1}\;\int\limits_{0}^{2}x_{1}\cdot x_{2}\;dx_{1}\;dx_{2}\;=\;1\) -- Always a good thing to check. Solve the first one. \(\int\limits_{0}^{1}\;\int\limits_{0}^{1}x_{1}\cdot x_{2}\;dx_{1}\;dx_{2}\) Solve the second one. \(\int\limits_{0}^{1}\;\int\limits_{0}^{x_{2}}x_{1}\cdot x_{2}\;dx_{1}\;dx_{2}\) Just look REALLY carefully at the limits. Tell if if the second one is just wrong!

OpenStudy (anonymous):

I get 1/8 for the second one

OpenStudy (anonymous):

and 1/ 4 for the first

OpenStudy (tkhunny):

Okay, then it is time to decide who it right. How should it work? What does the setup all mean? Have we wandered off, somewhere?

OpenStudy (anonymous):

I am sorry. I may have misunderstood your previous post. What are you referring to in the second integral? Solve the second one. ∫01∫0x2x1⋅x2dx1dx2 Just look REALLY carefully at the limits. Tell if if the second one is just wrong!

OpenStudy (tkhunny):

There are 3 integrals. What changed from the first (just check) to the second (answering question #1)?

OpenStudy (anonymous):

The bounds?

OpenStudy (tkhunny):

Right. Does this solve the problem? If so, how? If not, why not?

OpenStudy (anonymous):

Are you referring to the problem of me not knowing how to draw a picture of the region? If so, would the first part be the entire area = the entire probability from 0 to 1? Then the second part would take up approximately .25 of the area. The second part would take up 1/24 and from examining this geometrically, I should be able to come up with the marginal density?

OpenStudy (anonymous):

I am sorry if I seem really dim. But, I really appreciate the help

OpenStudy (tkhunny):

If the joint density were uniform (1/2 throughout). it would be as simply as drawing the picture and using the very square geometry. As it is NOT uniform, the geometry gives only a clue. This is why we need the integral. Solving the first: It is the VOLUME under the surface that is important. If we choose 0 < x1 < 1 and 0 < x2 < 1,clearly this is 1/2 the total area for a UNIFORM distribution. Upon integration with the actual joint density, we see that the volume in the first square (adjacent to the Origin) is only 1/4 - Leaving a whopping 3/4 in the other square. Solving the second: x1 * x2 is symmetrical about x1 = x2 in the first square., so this is easily answered by 1/2 of whatever we answered on the first question. x2 never exceeds Unity, so there can't be any in the 2nd square. I have to find exception with the given answer.

OpenStudy (anonymous):

you find exception with the given answer 1/24?